# PC-Gym: Benchmark Environments for Process Control Problems

### v2

@@ Abstract
PC-Gym is an open-source tool for developing and evaluating reinforcement learning (RL) algorithms in chemical process control. It features environments that simulate various chemical processes, incorporating nonlinear dynamics, disturbances, and constraints. The tool includes customizable constraint handling, disturbance generation, and reward function design. The tool enables comparison of RL algorithms against nonlinear Model Predictive Control (NMPC) across different scenarios. Case studies demonstrate its effectiveness in evaluating RL approaches for systems like continuously stirred tank reactors, multistage extraction processes, and crystallization reactors. The results reveal performance gaps between RL algorithms and NMPC oracles, highlighting areas for improvement. By providing a standardized platform, PC-Gym aims to accelerate research at the intersection of machine learning and process systems engineering. It connects theoretical RL advances with practical industrial process control applications, offering researchers a tool for exploring data-driven control solutions.
@@ 1 Introduction
Chemical processes exhibit complex dynamics and constraints that set them apart from traditional RL settings. These processes often involve multiple interacting variables, nonlinear behavior, and time-varying parameters, making them particularly difficult to control using standard RL techniques. Furthermore, chemical processes are often subject to strict safety, environmental, and economic constraints, which must be carefully considered in the control strategy. Violating these constraints can lead to severe consequences, such as equipment damage, product quality degradation, or even human harm.
@@ 1.1 Related Works
The development of reinforcement learning (RL) algorithms relies heavily on the availability of easily implementable environments that allow for efficient experimentation and iteration. Frameworks such as Gymnasium (formerly known as Gym) by Farama-Foundation [towers_gymnasium_2023] and OpenAI’s Gym [gym] have been instrumental in providing standardized discrete-time environments for RL research, offering a wide range of control problems, games, and simulations. However, these frameworks lack the latest parallelization methods for efficiently learning effective control policies and developing algorithms. To overcome this limitation, several libraries, such as Gymnax [gymnax2022github], Brax [brax2021github], Jumanji [bonnet2024jumanji], and PGX [koyamada2023pgx], have been developed to provide vectorized RL environments in JAX [jax2018github], enabling parallel simulations on Graphics Processing Units while maintaining the discrete-time nature of the environments. While these frameworks cover a variety of control problem types, there is still a need for domain-specific environments. Consequently, various natural sciences and engineering communities have started creating custom environments in the same format to simulate their specific problems, such as OR-Gym [HubbsOR-Gym] for operational research and ChemGymRL [chemgymrl] for chemical discovery. Despite these advancements, there remains a gap in the availability of process control environments with specific functionality tailored to creating discrete-time process control problems.
@@ 2.1 Dynamical systems
with stochasticity as the result of general uncertain parameters, $\textbf{w}_{t}\in\mathcal{W}\subseteq\mathbb{R}^{n_{w}}$, indicating disturbance or model parameters, with $\mathcal{W}=\text{supp}(p(\textbf{w}))$, being the support of its distribution. This work aims to characterize a set of transition distributions $p(\cdot)\in\mathcal{P}$ to represent meaningful process control problems. Each transition function is parameterized in general state-space form, with means to specify different disturbances as the user deems appropriate.
@@ 2.2 Learning-based Control
Given a dynamical system as defined in Equation 1, an initial state is sampled from a distribution $X_{0}\sim p_{0}(\textbf{x}_{0})$ where $p_{0}(\cdot)$ is an assumed initial state distribution. The control problem is defined by the control inputs $\textbf{u}_{t}$, which maximize a sequence of rewards, allocated by a function, $r_{t}:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$. Learning-based control aims to find a sequence of control inputs as a function of the current state based on data collected from a dynamical system over a fixed horizon $T$. The control problem is approximated with a fixed horizon to create an episodic learning approximation to the infinite horizon task at hand. Therefore, the objective is to find the optimal control policy $\pi^{*}:\mathcal{X}\rightarrow\mathcal{U}$ that maps state to control actions, maximizing the expected cumulative reward.
$\max_{\pi}\hskip 5.69054pt$ $\mathbb{E}_{\pi}\left[\sum^{T}_{t=0}\gamma^{t}r_{t}(X_{t},U_{t})\right]$
@@ 2.3 Reinforcement Learning
A popular method to solve the control problem described in the previous section is reinforcement learning (RL), where the major assumption applied is that the system is Markovian. This implies that the problem is modeled as a Markov Decision Process (MDP), and the system’s future is independent of the past, given the current state. The reinforcement learning framework (Figure 1) involves an agent and an environment, where the agent observes a state of the environment and inputs an action to the environment, which transitions it to a new state and returns a reward based on a hidden reward function.
@@
Multiple successful types of algorithms have emerged to find the optimal policy in the reinforcement learning framework. In the context of chemical process control, three advanced algorithms have shown particular promise: Proximal Policy Optimization (PPO), Soft Actor-Critic (SAC), and Deep Deterministic Policy Gradient (DDPG). These algorithms are well-suited to handle the continuous action spaces and complex dynamics often encountered in process control problems. Central to many RL algorithms is the Q-function, also known as the state-action value function. Mathematically, it can be expressed as the expectation of a discounted sum of random variables:
$Q(\textbf{x},\textbf{u})=\mathbb{E}_{\pi}\left[\sum_{t=t^{\prime}}^{T}\gamma^{t-t^{\prime}}r_{t}(X_{t},U_{t})\middle|X_{t^{\prime}}=\textbf{x},U_{t^{\prime}}=\textbf{u}\right],$ (2)
@@
This formulation explicitly shows the Q-function as the expected sum of discounted future rewards, starting from state x and taking action u, then following the optimal policy for all subsequent steps. Another key concept is the value function $V(\textbf{x})$, which represents the expected return starting from state x and following the policy $\pi$ thereafter. It can be defined as,
$V^{\pi}(\textbf{x})=\mathbb{E}_{\pi}\left[\sum_{t=t^{\prime}}^{T}\gamma^{t-t^{\prime}}r_{t}({X}_{t},U_{t})\middle|X_{t^{\prime}}=\textbf{x}\right],$ (3)
@@
where $\gamma\in[0,1]$ is the discount factor. The value function helps in evaluating how “good” it is to be in a particular state, considering the long-term rewards that can be obtained from that state under the current policy. The advantage function, which is used in many RL algorithms, combines the Q-function and the value function:
$\hat{A}_{t}=Q(\textbf{x}_{t},\textbf{u}_{t})-V(\textbf{x}_{t})\,.$ (4)
@@
This function estimates the improvement in performance by taking a given control $\textbf{u}_{t}$, in state $\textbf{x}_{t}$, and following the policy thereafter, rather than simply following the policy. This provides an effective quantity for approximate policy improvement steps.
@@
In practice, these functions are typically parameterized using neural networks due to their ability to handle continuous state and action spaces. The Q-function can be parameterized by parameters $\theta$ as $Q_{\theta}(\textbf{x}_{t},\textbf{u}_{t})$, and the policy as $\pi_{\phi}(\textbf{u}_{t}|\textbf{x}_{t})$ with parameters $\phi$. The parameterized policy can be either deterministic, directly outputting actions $\textbf{u}_{t}=\pi_{\phi}(\textbf{x}_{t})$, or stochastic, outputting a probability distribution over actions. For continuous control problems in chemical processes, the policy network often outputs the mean and standard deviation of a Gaussian distribution:
$\pi_{\phi}(\textbf{u}_{t}|\textbf{x}_{t})=\mathcal{N}(\boldsymbol{\mu}_{\phi}(\textbf{x}_{t}),\boldsymbol{\sigma}_{\phi}(\textbf{x}_{t})),$ (5)
where $\boldsymbol{\mu}_{\phi}(\textbf{x}_{t})$ and $\boldsymbol{\sigma}_{\phi}(\textbf{x}_{t})$ are neural network outputs. The value function is similarly parameterized as $V_{\psi}(\textbf{x}_{t})$ with parameters $\psi$. These parameterizations transform the reinforcement learning problem into one of finding optimal parameters $\theta$, $\phi$, and $\psi$ through gradient-based optimization. The neural networks are typically trained by minimizing temporal difference errors for the value functions and maximizing expected returns for the policy, subject to various constraints that depend on the specific algorithm employed. In the following, we give a brief description of three learning algorithms that have proved effective policy optimizers within the literature.
@@ 2.3.1 Soft Actor-Critic (SAC)
SAC is an off-policy actor-critic method that incorporates entropy regularization into the objective function [haarnoja2018soft]. The algorithm employs experience replay and uses target networks for both critics with soft updates. It maintains two Q-function networks and one policy network, all implemented as fully connected neural networks. It aims to learn a policy that maximizes both the expected return and the entropy of the policy. The objective function $J(\pi)$ represents the expected return of the policy and is defined as:
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}(X_{t},U_{t})\right]\,.$ (6)
SAC then augments this objective with an entropy term:
$J_{SAC}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}(X_{t},U_{t})+\alpha H(\pi)\right]\,,$ (7)
where $H(\pi)$ is the entropy of the policy, and $\alpha$ is a temperature parameter that automatically adjusts through gradient descent to maintain a target entropy level. SAC learns a stochastic policy by outputting the parameters of a Gaussian distribution, which can be beneficial in chemical process environments with multiple optimal solutions or where exploration is crucial.
@@ 2.3.2 Deep Deterministic Policy Gradient (DDPG)
DDPG is an off-policy actor-critic algorithm designed specifically for continuous action spaces, making it particularly suitable for process control tasks [DDPG]. For complete implementation details, readers are referred to the original paper. The algorithm implements experience replay and employs target networks for both actor and critic with soft updates. The networks are fully connected, where the actor outputs deterministic actions and the critic network estimates the Q-value of state-action pairs. The actor is updated using the deterministic policy gradient:
$\nabla_{\phi}J=\mathbb{E}_{\textbf{x}_{t}\sim\rho^{\beta}}\left[\nabla_{a}Q(\textbf{x},\textbf{u};\theta)\mid_{{\textbf{x}=\textbf{x}_{t},\textbf{u}=\mu(\textbf{x}_{t};\phi)}}\nabla_{\phi}\boldsymbol{\mu}(\textbf{x};\phi)\mid_{{\textbf{x}=\textbf{x}_{t}}}\right]$ (8)
where $\boldsymbol{\mu}(\textbf{x};\phi)$ is the deterministic policy and $Q(\textbf{x},\textbf{u};\theta)$ is the critic’s Q-function. The term $\rho^{\beta}$ denotes the state distribution under the behavior policy $\beta$. This behavior policy adds Ornstein-Uhlenbeck noise to the deterministic actions for exploration. The state distribution $\rho^{\beta}$ represents the probability distribution of states encountered when following this behavior policy.
@@ 2.3.3 Proximal Policy Optimization (PPO)
PPO is an on-policy policy gradient method that aims to improve the stability and sample efficiency of policy updates [schulman2017proximalpolicyoptimizationalgorithms]. Unlike SAC and DDPG, PPO does not use experience replay or target networks, instead collecting trajectories using the current policy for updates. It uses an actor-network that outputs the mean and standard deviation of a Gaussian policy and a critic network that estimates the value function. PPO introduces a clipped surrogate objective function:
$L^{CLIP}(\theta_{k})=\min\left(\frac{\pi_{\theta_{k}}}{\pi_{\theta_{k-1}}}\hat{A}_{t},\text{clip}\left(\frac{\pi_{\theta_{k}}}{\pi_{\theta_{k-1}}},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\,,$ (9)
where $\frac{\pi_{\theta_{k}}}{\pi_{\theta_{k-1}}}$ is the probability ratio between the new and old policies, $\hat{A}_{t}$ is the estimated advantage function calculated using Generalized Advantage Estimation (GAE), and $\epsilon$ is a hyperparameter that controls the clipping range. This objective function is a surrogate to Equation 6, which imposes an approximate trust-region and ensures that the policy updates are neither too large nor too small, leading to more stable learning. This is particularly beneficial for highly nonlinear chemical processes.
Throughout this paper, we will use PPO, SAC, and DDPG to demonstrate the flexibility of PC-Gym. These algorithms have been chosen for their effectiveness in dealing with continuous action spaces and their ability to handle the complexities often encountered in chemical process control problems. By using PC-Gym, we aim to provide insights into the relative strengths and weaknesses of these algorithms across different process control scenarios.

### v4

@@ Abstract
PC-Gym is an open-source tool for developing and evaluating reinforcement learning (RL) algorithms in chemical process control. It features environments that simulate various chemical processes, incorporating nonlinear dynamics, disturbances, and constraints. The tool includes customizable constraint handling, disturbance generation, reward function design, and enables comparison of RL algorithms against Nonlinear Model Predictive Control (NMPC) across different scenarios. Case studies demonstrate the framework’s effectiveness in evaluating RL approaches for systems like continuously stirred tank reactors, multistage extraction processes, and crystallization reactors. The results reveal performance gaps between RL algorithms and NMPC oracles, highlighting areas for improvement and enabling benchmarking. By providing a standardized platform, PC-Gym aims to accelerate research at the intersection of machine learning, control, and process systems engineering. By connecting theoretical RL advances with practical industrial process control applications, offering researchers a tool for exploring data-driven control solutions.
@@ 1 Introduction
Chemical processes exhibit dynamics and constraints that set them apart from traditional RL settings. These processes often involve multiple interacting variables, nonlinear behavior, and time-varying parameters, making them particularly difficult to control using standard RL techniques [nian_review_2020, shin_reinforcement_2019]. Furthermore, chemical processes are often subject to strict safety, environmental, and economic constraints, which must be carefully considered in the control strategy. Violating these constraints can lead to severe consequences, such as equipment damage, product quality degradation, or even human harm.
@@ 1.1 Related Works
[…]
@@ 2.1 Dynamical systems
with stochasticity as the result of general uncertain parameters, $\textbf{w}_{t}\in\mathcal{W}\subseteq\mathbb{R}^{n_{w}}$, indicating disturbance or model parameters, with $\mathcal{W}=\text{supp}(p(\textbf{w}))$, being the support of its distribution. This work aims to characterize a set of transition density functions $p(\cdot)\in\mathcal{P}$ to represent meaningful process control problems. Each transition function is parameterized in general state-space form, with means to specify different disturbances as the user deems appropriate.
@@ 2.2 Learning-based Control
Given a dynamical system as defined in Equation 1, an initial state is sampled from a distribution $X_{0}\sim p_{0}(\textbf{x}_{0})$ where $p_{0}(\cdot)$ is an assumed initial state distribution. The control problem is defined by the control inputs $\textbf{u}_{t}$, which maximize a sequence of rewards, allocated by a function, $r_{t+1}:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$. Learning-based control aims to identify a state feedback policy based on data collected from a dynamical system over a fixed horizon $T$. Here, the control problem is approximated with a fixed horizon to create an episodic learning approximation to the infinite horizon task at hand. Therefore, the objective is to find the optimal control policy $\pi^{*}:\mathcal{X}\rightarrow\mathcal{U}$ that maps state to control actions, maximizing the expected cumulative reward.
$\max_{\pi}\hskip 5.69054pt$ $\mathbb{E}_{\pi}\left[\sum^{T-1}_{t=0}\gamma^{t}r_{t+1}(X_{t},U_{t})\right]$
@@ 2.3 Reinforcement Learning
A popular method to solve the control problem described in the previous section is reinforcement learning (RL), where the major assumption applied is that the system is Markovian and that the problem is modeled as a Markov Decision Process (MDP). Figure 1 describes the MDP formulation as the interaction between a control policy (agent) and a discrete-time stochastic decision process (environment). The control policy observes the state of the process and inputs a control decision to the environment, the environment then transitions to a new state and returns a reward based on a reward function.
@@
Many types of algorithms have emerged to find an approximate optimal policy in the RL framework. In the context of chemical process control, three advanced algorithms have shown particular promise: Proximal Policy Optimization (PPO), Soft Actor-Critic (SAC), and Deep Deterministic Policy Gradient (DDPG). These algorithms are well-suited to handle continuous action spaces and complex dynamics often encountered in process control problems. Central to many RL algorithms is the Q-function, also known as the state-action value function. Mathematically, it can be expressed as the expectation of a discounted sum of scalar random variables:
$Q^{\pi}(\textbf{x}_{t},\textbf{u}_{t})=\mathbb{E}_{\pi}\left[\sum_{k=t}^{T-1}\gamma^{k-t}r_{k+1}(X_{k},U_{k})\middle|X_{t}=\textbf{x}_{t},U_{t}=\textbf{u}_{t}\right]\,.$ (2)
@@
This formulation explicitly shows the Q-function as the expected sum of discounted future rewards, starting from state $\textbf{x}_{t}$ and taking control $\textbf{u}_{t}$, then following the policy for all subsequent steps. Another key concept is the value function $V^{\pi}(\textbf{x}_{t})$, which represents the expected return starting from state $\textbf{x}_{t}$ and following the policy $\pi$. It can be defined as,
$V^{\pi}(\textbf{x}_{t})=\mathbb{E}_{\pi}\left[\sum_{k=t}^{T-1}\gamma^{k-t}r_{k+1}({X}_{k},U_{k})\middle|X_{t}=\textbf{x}_{t}\right],$ (3)
@@
where $\gamma\in[0,1]$ is the discount factor and $U_{k}$ represents the control action generated by policy $\pi$ at each timestep, $k\in\{t,\ldots,T-1\}$. The value function helps in evaluating how “good” it is to be in a particular state, considering the long-term rewards that can be obtained from that state under the current policy. The advantage function, which is used in many RL algorithms, combines the Q-function and the value function:
$A^{\pi}(\textbf{x}_{t},\textbf{u}_{t})=Q^{\pi}(\textbf{x}_{t},\textbf{u}_{t})-V^{\pi}(\textbf{x}_{t})\,.$ (4)
@@
This function estimates the difference in performance between taking a given control $\textbf{u}_{t}$, in state $\textbf{x}_{t}$, and following the policy thereafter, rather than simply following the policy. This provides an effective quantity for approximate policy improvement steps.
@@
In practice, these functions are typically parameterized using neural networks due to their ability to handle continuous state and control spaces. The Q-function can be parameterized by $\theta$ as $Q_{\theta}(\textbf{x}_{t},\textbf{u}_{t})$, and the policy as $\pi_{\phi}(\textbf{u}_{t}|\textbf{x}_{t})$ with parameters $\phi$. The parameterized policy can be either deterministic, directly outputting controls $\textbf{u}_{t}=\pi_{\phi}(\textbf{x}_{t})$, or stochastic, outputting a probability density over controls. For continuous control problems in chemical processes, the policy network often outputs the mean and standard deviation of a Gaussian distribution:
$\pi_{\phi}(\textbf{u}_{t}|\textbf{x}_{t})=\mathcal{N}(\boldsymbol{\mu}_{\phi}(\textbf{x}_{t}),diag(\boldsymbol{\sigma}_{\phi}(\textbf{x}_{t}))),$ (5)
where $\boldsymbol{\mu}_{\phi}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{u}}$ and $\boldsymbol{\sigma}_{\phi}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{u}}$ are neural network outputs [^2] It is worth noting in deployment, Gaussian policies may be transformed into deterministic controllers by selecting the control corresponding to the mean. Control bounds may be enforced by a hyperbolic tangent function [haarnoja2018soft] or simply clipping the policy output.. The value function is similarly parameterized as $V_{\psi}(\textbf{x}_{t})$ with parameters $\psi$. These parameterizations transform the reinforcement learning problem into one of finding optimal parameters $\phi$, $\theta$ and $\psi$. These are generally estimated through gradient-based optimization with gradients estimated through the policy gradient theorem [sutton_policy_nodate] in the case of $\phi$, and variations of temporal difference (TD) learning in the case of $\theta$ [watkins1992q] and $\psi$ [boyan1999least]. The policy gradient provides a direction in the parameter space to maximize the expected returns for the policy. TD learning provides a means to identify a bootstrapped estimate of the squared error associated with a Q-function or value function approximation under a given behavior [suttonRL] or target policy [watkins1992q]. Taking the negative of the gradient with respect to this estimate provides a direction in the parameter space to improve the approximation. In the following, we give a brief description of three learning algorithms that have proved effective policy optimizers within the literature.
@@ 2.3.1 Soft Actor-Critic (SAC)
SAC is an off-policy actor-critic method that incorporates entropy regularization into the objective function [haarnoja2018soft]. Actor-critic methods parameterize both a policy (actor) and a Q-function or a value function (critic). The SAC algorithm employs both a value function and a Q-function together with a policy. Additionally, it uses experience replay and target networks [mnih_human-level_2015] for both the Q-function and value function approximation, which are updated with soft updates. It aims to learn a policy that maximizes both the expected return and the entropy of the policy. In an episodic setting, the objective function $J(\pi)$ represents the expected return of the policy and is defined as:
$J(\pi_{\phi})=\mathbb{E}_{\pi_{\phi}}\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t+1}(X_{t},U_{t})\right]\,.$ (6)
SAC augments this objective with an entropy term:
$J_{SAC}(\pi_{\phi})=J(\pi_{\phi})+\alpha H(\pi_{\phi})\,,$ (7)
where $H(\pi_{\phi})$ is the expected entropy of the policy given the associated state distribution, and $\alpha$ is a temperature parameter that can be automatically adjusted through gradient descent to maintain a target entropy level. SAC learns a stochastic policy that parameterizes the moments of a Gaussian distribution conditional to the current state, which can be beneficial in chemical process environments with multiple optimal solutions or where exploration is crucial [pmlr-v97-ahmed19a].
@@ 2.3.2 Deep Deterministic Policy Gradient (DDPG)
DDPG is an off-policy actor-critic algorithm designed specifically for continuous control spaces, making it particularly suitable for process control tasks [DDPG]. For complete implementation details, readers are referred to the original paper. The algorithm implements experience replay and employs target networks for both policy and Q-function with soft updates. The networks are fully connected, where the policy outputs deterministic controls. The actor is updated using the deterministic policy gradient:
$\nabla_{\phi}J=\mathbb{E}_{\textbf{x}_{t}\sim\rho^{\beta}}\left[\nabla_{\textbf{u}}Q(\textbf{x},\textbf{u};\theta)\mid_{{\textbf{x}=\textbf{x}_{t},\textbf{u}=\mu(\textbf{x}_{t};\phi)}}\nabla_{\phi}\boldsymbol{\mu}(\textbf{x};\phi)\mid_{{\textbf{x}=\textbf{x}_{t}}}\right]$ (8)
where $\boldsymbol{\mu}(\textbf{x};\phi)$ is the deterministic policy and $Q(\textbf{x},\textbf{u};\theta)$ is the Q-function. The term $\rho^{\beta}$ denotes the state distribution under the behavior policy $\beta$. This behavior policy adds Ornstein-Uhlenbeck noise to the deterministic controls for exploration. The state distribution $\rho^{\beta}$ represents the probability distribution of states encountered when following this behavior policy.
@@ 2.3.3 Proximal Policy Optimization (PPO)
PPO builds on Trust Region Policy Optimization (TRPO) [schulman_trust_2015], presenting a more computationally efficient approach to stable policy updates [schulman2017proximalpolicyoptimizationalgorithms]. While TRPO enforces a trust region through a KL-divergence constraint, PPO achieves similar stability through a simpler clipped objective function. Unlike SAC and DDPG, PPO does not use experience replay or target networks, instead collecting trajectories using the current policy for updates. It uses a policy-network that outputs the mean and standard deviation of a Gaussian policy and a value function network and introduces a clipped surrogate objective function:
$L^{CLIP}(\phi_{j},\phi_{j-1})=\mathbb{E}_{(\textbf{x},\textbf{u})\sim\pi_{\phi_{j-1}}}\left[\min\left(\frac{\pi_{\phi_{j}}}{\pi_{\phi_{j-1}}}A_{\psi}(\textbf{x},\textbf{u}),\text{clip}\left(\frac{\pi_{\phi_{j}}}{\pi_{\phi_{j-1}}},1-\epsilon,1+\epsilon\right)A_{\psi}(\textbf{x},\textbf{u})\right)\right]\,,$ (9)
where $\frac{\pi_{\phi_{j}}}{\pi_{\phi_{j-1}}}$ is the probability ratio between the new and old policies, $A_{\psi}(\textbf{x},\textbf{u})$ is the estimated advantage function calculated using Generalized Advantage Estimation (GAE) [^3] Although the definition provided by Eq. 4 indicates that the advantage estimate would require parameterization of both the state-action value function and the value function, one only needs the latter. Please see the cited work [schulman_high-dimensional_2018] for more information. [schulman_high-dimensional_2018], $j$ is the policy update step and $\epsilon$ is a hyperparameter that controls the clipping range. This objective function is a surrogate to Equation 6, which imposes an approximate trust-region and ensures that the policy updates are neither too large nor too small, leading to more stable learning. This is particularly beneficial for highly nonlinear chemical processes.
Throughout this paper, we will use PPO, SAC, and DDPG to demonstrate the capabilities of PC-Gym. These algorithms have been chosen for their effectiveness in dealing with continuous control spaces and their ability to handle the complexities often encountered in chemical process control problems. By using PC-Gym, we aim to provide insights into the relative strengths and weaknesses of these algorithms across different process control scenarios.